PLEASE HELP ASAP: Prof Davis counts his midterm as 2/3 of the grade and his final as 1/3 of the grade. Jason scored only 64 on the midterm. What range of scores on the final exam would put Jason's average between 70 and 79 inclusive? Solve by using a compound inequality. PLEASE HELP
thank you

Answer by checkley71(8403)

(2/3)64+(1/3)X=70 TO 79
128/3+X/3=70 TO 79
(128+X)/3=70 TO 79
128+X=210 OR 237
X=210-128 OR 82 & X=237-128 OR X=109 THEREFORE TO AVERAGE BETWEEN 70 & 79
THE FINAL TEST SCORE NEEDS TO BE BETWEEN 82 & 100.